The revenue R received by a company selling x pairs of sunglasses per week is given by the function R(x)=−0.1x2+80x.

​(a) Find the values of R(13) and R(35).

(b) How many pairs of sunglasses must be sold in order for the revenue to be $ 7000 per​ week?

(c) How many pairs of sunglasses must be sold in order for revenue to be $16,000 per​ week?
Answer #1

R2 = -0.1 202 +80k (a) R (13) = -0.18/13) ~ + 80x13= 1023.1 cho ..R(13) = 1023.1 R (35) = – 0-1X35 + 80x 35 yr (35) = 2677-5